# Simple Hamiltonian dynamics is a powerful quantum processing resource

### v1

@@ V Entanglement and Integrability
Given the above performance benchmarking using the various quantum reservoir models, we now turn our attention to the properties of these quantum reservoirs. Entanglement is known to be necessary for universal quantum computation and is often discussed in quantum machine learning [56]. In the benchmark, all the quantum reservoirs we evaluated can in general achieve a significantly better performance in a comparison to the performance random single qubit rotations only can provide. All the quantum reservoirs have the power to generate entanglement. As we have seen in quantum computation, entanglement has a certain role in the performance of QERC. However, the reverse statement does not hold, that is, the ZZ-Ising model with $g/J_{0}=0$ does not give any improvement. This may not so surprising, however the comparison of this model with the ZZ-X model gives us an interesting insight. The difference between the ZZ-Ising model with $g/J_{0}=0$ and the ZZ-X model is the last layer of single-qubit gates in the circuit for the ZZ-X model, yet the ZZ-X model can achieve as high as any other quantum reservoir models. Fig. 5 (a) shows the performance (accuracy rate against the single qubit rotation $\theta_{x}$. The optimal points appear around the angle $\pi/8$, which is equivalent to the T-gate. As we know that single qubit rotations alone cannot contribute to the performance, this indicates that the changing the measurement axis is enough to extract of the entangling dynamics the ZZ-Ising interaction created for the computation. This suggests that the complexity the ZZ-Ising model can generate is sufficient for high QERC performance, which agrees with the XX-Ising models performance noting that the XX and ZZ interactions are related by a simple unitary transformation. A similar behaviour can be seen with the Clifford circuit and the Clifford +T model. This increase of the computational complexity has been reported with Clifford circuit with a single T gate [57], however in our case a single T gate is not enough to fully recover the accuracy rate.
@@
Next, we look into the influence of the integrability of the XX-Ising model on its performance as a quantum reservoir. In the benchmark, we set $\alpha$ to be 1.5, as that parameter value gives the optimal performance. To consider integrability in this case, we take different values of $\alpha$ to determine how its integrability affects the performance. We will consider three different interaction ranges ($\alpha=0.0,1.5$ and $\infty$). The first $\alpha=0$ corresponds to all-to-all connectivity with equal strength, whereas $\alpha=\infty$ corresponds to nearest neighbour coupling only. Now $\alpha=1.5$ sits between this two extremes and corresponds to the long range order parameter present in the recent ion-trap experiments [45, 8, 47, 48]. The training accuracy rates are summarized for these three parameter values in Fig. 6.
@@
Figure 6: Plot of the training accuracy rate of XX-Ising model for three different values of $\alpha=0.0$(blue), $1.5$(green), $\infty$(red). Here we have set $g/J_{0}=0.0$. The shadows displaced show the standard deviations obtained from 40 to 50 epochs. The inset table shows the unitary operators at the stroboscopic times, $J_{0}t=n\pi,(n+1/4)\pi,(n+1/2)\pi$ and $(n+1/3)\pi$ with $n\in\mathbb{N}_{\geq 0}$ and where $N$ is the number of sites and $\hat{P}_{X}$ the parity operator along the x-axis.
@@
The best accuracy can be achieved for a middle range $\alpha$ value. Unlike the case of $\alpha=1.5$, the $\alpha=0$ and $\infty$ cases show periodical behaviour. This reflects to the periodic dynamics of the Hamiltonian with these particular values. It is useful to explore such and mechanics more analytically. Our unitary operator can be decomposed as
@@
where the XX-interactions between each site commute with one another. In this case the total unitary can be decomposed as a product of the simple oscillators with the angular frequencies $\omega_{lm}={J_{0}}/{|l-m|^{\alpha}}$. Thus, when $\alpha=1.5$, the oscillators are in different angular frequencies. This means the recurrence of dynamics is avoided, resulting in non-periodical dynamics.
Now in the nearest-neighbour interaction case ($\alpha=\infty$), the oscillators may share the same angular frequency introducing a symmetry. Our unitary operator may then be written as
@@
It is interesting to also explore the opposite limit of $\alpha=0$. Here the Ising model also recovers the system symmetry, and the dynamics becomes solvable as
@@
with all oscillators having the same angular frequency $\omega_{lm}=J_{0}$. Thus, similar to $\alpha=\infty$, the dynamics is analytic at the times $J_{0}t=n\pi,(n+1/4)\pi,(n+1/2)\pi,(n+3/4)\pi$, where its performance reduces, as shown in Fig. 6. When the dynamics is periodical, unless the period is extremely long, it is likely that the dynamics can be fully supported by a small subspace of the Hilbert space. As shown in Fig. 4, the performance quickly recovers as $\alpha$ increases from zero and peaks around $\alpha=1$, and then gradually decreases to the large $\alpha$ limit. This indicates that the non-integrability of the quantum system contributes to the performance of the quantum neural networks. Interestingly, the power to generate entanglement is not sufficient to estimate the QERC performance as we have seen in above with $\hat{U}^{(XX)}_{\alpha=\infty}$ at $J_{0}t=(n+1/4)\pi$ for instance.
Integrability impacts the performance of the QERC to some degree, however it does not give us clear criteria for the quantum reservoir. The ZZ-X model is an integrable system thought it does not hold U(1) symmetry as a result of its time evolution which consists of two Hamiltonian components. This indicates that when the time-evolution has a “gate” structure the integrability may not work as an indicator. This is similar to the case of the Clifford gates $+$ T-gates. The Clifford gate circuit can be efficiently simulated, and does not give us high performance as a quantum reservoir. However the one-depth of T-gates can be enough to achieve similar high performance to the Haar measure sampling. It may not be surprising that the integrability of the Hamiltonian system can significantly affect the performance as quantum neural network. However, the integrability is much less important in the ZZ-X model, which is integrable with the analytical form for the output state. Integrability and periodicity are somewhat related in classical dynamics, and as we have seen above, the periodicity in the Hamiltonian dynamics prevents the states from evolving sufficiently for the one-layer classical neural network to distinguish their distributions. Non-integrability and non-periodicity of the system dynamics are not a necessary condition for a good quantum reservoir, however they can still serve as an indication to provide good neural networks as well as CUE does, which can be useful in designing quantum reservoirs.
@@ VI Information scrambling and sampling cost
In the previous section, we analyzed the properties of the quantum reservoirs with respect to their complex dynamics. Here, we focus on the other element of quantum computation that can be achieved by quantum reservoir, that is, the convergence to the solution through the quantum algorithm. The quantum dynamics provides a feature space for this computational model, and hence the dynamics complexity is important to deliver information scrambling necessary. To reach a solution, the conversion to it needs to happen, and it is not clear if the quantum reservoir carries the role of information scrambling only or not. If the quantum reservoir somehow contributes to the convergency process, the cost of measurement at M-layer would improve dramatically. To investigate this, we need to estimate the sampling cost for the measurement.
@@
The amplitude distribution of the final state of the quantum reservoir was used as the output of the M-layer in our benchmark, however in the simulations we used the amplitude distribution reconstructed from the data. Fig. 7 shows the behaviour of the accuracy rates dependent on the sampling number. All the models in this comparison achieve similar accuracy rates with the theoretical amplitude distribution when the sampling number is large enough. To directly see how well the distribution is reconstructed, we employ the Kullback-Leibler divergence (KLD). Please see Appendix B.2 for more details. In the comparison of these two figures in the Appendix B.2 , the accuracy rate converges to the theoretical value before the distributions do. The necessary sampling numbers in the implementation of the QERC are not high in comparison to the dimension of the Hilbert space. In addition, while the cost of the statistical reconstruction of the distribution function is dependent on the size of the computational basis, the necessary sampling number seems less dependent on the size of the computational basis.
Whereas the behaviour of KLD for the different quantum reservoirs remains the same, the ZZ-X model exhibits a different tendency in the small sampling number regime, where it requires a much smaller sampling number than the other models require to achieve the same accuracy (see Appendix B.2). This can not only be a large advantage in its implementation. but also be a hint to consider the trade-off between the information scrambling and the convergence to the classes.
Information scrambling has to happen during the time evaluation of the quantum reservoir. If this is only the role the quantum reservoir carries, the CUE should give the best result. In QERC, the amplitude distribution the M-layer generates has to be fed into the ONN, so if the ONN feels the task to categorize these distribution difficult, the QERC would not achieve a high accuracy rate. The advantage of the ZZ-X model for the sampling number indicates that the distributions the ZZ-X model generates are easier to categorize for the ONN. To examine this ansatz, we investigate the participation ratio in the next section.
@@ VII Participation Ratio (PR)
To evaluate the power of reservoir computation, the Lyapunov exponent and memory capacities such as $\tau$-capacity are widely used: the former being a measure of the chaotic dynamics and nonlinearlity in classical reservoirs [58], while the latter are used in classical and quantum reservoir computing [59, 30, 31, 60, 61, 62]. The memory capacity is to measure how much past information is stored in the reservoir, and unfortunately such a measure is not appropriate to evaluate the quantum reservoirs used in QERC. Also, to see what is happening in the computational process though the QERC, we need to know what complexity is generated in the state as well as its degree. Hence we employ the participation ratio (PR) used in condensed matter physics to measure the local and extended states [63], such as the Anderson localized and ergodic states, to evaluate the dynamics. For a quantum state $|\Psi\rangle=\sum_{l}c_{l}|l\rangle$, the PR is defined as,
@@
The PR of the output state can be highly dependent on the initial state, and so we considered the averaged PR (APR) over the input state set $\mathbb{D}$ defined by,
@@
This value tells us how widely the state is spread over the computational basis on average.
Now if the APR is high, an input state would evolve to an output state supported by many computational basis states. Hence it is more likely that the dynamics yields a strong information scrambling. However a set of such states could be more difficult to categorize by the ONN. If the distribution function reconstructed from the sampling has a distinct feature for each class of the classification task, it would be easy for the ONN to classify these reconstructed distributions, and hence, the sampling number required would be small. To capture such the behavior of all input states contributions, we consider the PR of the mixed states over all inputs, reprented by,
@@
Then, its PR is obtained by,
${\mathrm{IAPR}}\left(\mathbb{D}\right)=\frac{{(\#\,\mathrm{states})}^{2}}{{\displaystyle\sum_{l}\left({\sum_{|\Psi\rangle\in\mathbb{D}}|\langle l|\Psi\rangle|^{2}}\right)^{2}}},$ (12)
and we refer to it as the improved APR (IAPR).
Remembering that $\mathrm{APR}\leq\mathrm{IAPR}$, our analysis considers the difference between these defined by $\Delta\mathrm{PR}=\mathrm{IAPR}-\mathrm{APR}$. If the APR is large, IAPR must be large, however when the APR is small, the IAPR can be range from small to large. If the reservoir dynamics creates distributions with a large IAPR but small APR, it might be the optimal output for the ONN to classify. Hence, for our purpose, the measure of our interest is $\Delta\mathrm{PR}$.
@@
Before analysing the PRs, we need to take the issue of the bias within the input state set into account. These PR values are dependent on the trend and bias within the set of the input states. In order to evaluate the quantum reservoir neutral to these treads, we introduce two random input sets. First, we consider uniformly distributed points in the $2^{N}$-dimensional space. We generate $10,000$ points randomly and use the same encoder as we did in the benchmark (described in Ref. [21] and the Appendix A.1). The completely random input set is however not so realistic, as the input data set generally has some bias. Hence we also consider a random data set generated by isotropic Gaussian blobs in the high dimensional square with length $-10$ to $10$ (see the sketch in Fig. 8 (b-1)). Here we randomly generate $10,000$ points in the high dimensional space. These points form $10$ isotropic Gaussian blobs (clusters) where the standard deviation of the clusters is $1.2$. These points are mapped to the angles following the encoding process mentioned in Ref. [21] (and Appendix A.1). We also considered the MNIST data set to evaluate the actual practical case (see (c-1) in Fig. 8).
@@
Figure 8: Plot of the APR vs. $\Delta\mathrm{PR}$ for ZZX-Ising model using ten qubits. The red and green gradation points show the result with different transverse magnetic strengths $0\leq g/J_{0}\leq 5$, where we used normal and randomized encoders in red and green, respectively. Here, we set $J_{0}t=3.5$ with $\alpha=1.5$. The blue dots show the change of APR and IAPR thought the time evolution. The red and blue triangle markers show the results of the Haar and SRC reservoirs. Now (a) and (b) show the results of sets where $10,000$ initial states were randomly generated. For (a), all states were created by angles $\theta_{l}$ and $\phi_{l}$ that are uniformly randomly generated. Next (b) illustrates the result for the initial state set generated by isotopic Gaussian blobs while (c) depicts the results for the MNIST dataset using 60,000 images.
@@
We show the the APR and $\Delta\mathrm{PR}$ values for the ZZ-Ising model with the benchmark against two random reservoirs: the Haar measure sampling and the single rotation random circuit (SRC). These are given by the red and blue triangles in Fig. 8 respectively. For the ZZ-Ising situation thedifference between the randomization of the input state appears in the difference between the red (no randomization) and green (randomization) dots. As the magnetic field strength $g$ increases both the APR and IAPR decrease as shown with the red dots. This agrees with the lower accuracy seen without the randomization for the high $g$ regime. For all of the three data sets, the ZZ-Ising model get close to the value given by the Haar sampling, and hence by taking an optimal parameter values the behaviour of this reservoir is expected to be similar to that of Haar sampling. However, the optimal value for $g$ is not where the APR is the largest. To see this more clearly, we need to investigate the PR values for the ZZ-X model.
@@
The PR values for the ZZ-X model is shown in Fig. 9 with an interesting characteristic being that there are points above the Haar sampling in terms of the APR in the case of the ZZ-X model. Even more interestingly, the optimal point for the QERC performance does not appear near the Haar sampling. The blue dots in these figures show the dependency of the PRs on the single-rotation angle. The optimal point for $\theta_{x}$ in Fig. 5 (c) are near $\pi/8$ meaning the optimal point for this model clearly differs from the point given by Haar sampling. Further the ZZ-X model optimal point is characterized by a lower APR and higher IAPR. For the cases with a lower APR and a higher IAPR, each output distribution function does not broadly spread on the computational basis. This supports the ideal that the ZZ-X model utilize the Hilbert space as much as other models to support the set of the output states however the distribution of each output state is not so flat and hence it is easier for the ONN to distinguish them. This PR ratio and the high performance of the ZZ-X model provides evidence that the ZZ-X model realizes a partial convergence to the classification of the output distributions.

### v2

@@ V Entanglement and Integrability
Given the above performance benchmarking using the various quantum reservoir models, we now turn our attention to the properties of these quantum reservoirs. Entanglement is known to be necessary for universal quantum computation and is often discussed in quantum machine learning [59]. In the benchmark, we observe that the quantum reservoirs that have achieved a higher accuracy than the classical linear classifier provide generate entanglement in the computational process. By contrast, the quantum reservoir of random single-qubit rotations gives performance poorer than the classical linear classifier. This implies that entanglement is a key resource for the performance of QERC. Within the numerical observation, QERC is not an exception to quantum algorithms in general regarding the role of entanglement [60]. It is also known that entanglement does not guarantee the computational power that quantum computation promises. This is also shown in our benchmark, which we will explore next.
@@
The ZZ-Ising model with $g/J_{0}=0$ does not give any improvement, although it generates entanglement in the output state. This arises simply from the symmetry shared by both the dynamics and the measurement. The measurement used at the M-layer does not access to the effect of the entanglement generated in the state. Hence, the comparison of this model with the ZZ-X model gives us an interesting insight. The difference between the ZZ-Ising model with $g/J_{0}=0$ and the ZZ-X model is the last layer of single-qubit gates in the circuit for the ZZ-X model, yet the ZZ-X model can achieve as high as any other quantum reservoir models. Fig. 4 (a) shows the performance (accuracy rate against the single qubit rotation $\theta_{x}$. The optimal points appear around the angle $\pi/8$, which is equivalent to the T-gate. As we know, single-qubit rotations alone cannot contribute to the performance; this indicates that changing the measurement axis is sufficient to extract the entangling dynamics of the ZZ-Ising interaction created for the computation. This suggests that the complexity the ZZ-Ising model can generate is sufficient for high QERC performance, which agrees with the XX-Ising model performance, noting that the XX and ZZ interactions are related by a simple unitary transformation. A more interesting comparison can be seen with the Clifford circuit and the Clifford +T model. It is known that a Clifford circuit can be efficiently simulated in a classical computation [61, 62] despite of the entanglement in the computational system. Our numerical simulation shows that the single layer of T gates increases the accuracy rate. The role of the T-gates is similar to the computational complexity in learning a probability distribution, where a single T-gate was used [63]. In our case, a single T gate is insufficient to recover the accuracy rate fully.
@@
Next, we look into the influence of the integrability of the XX-Ising model on its performance as a quantum reservoir. Here, we consider the analytical form of the unitary map of the quantum reservoir and refer to its integrability. In the benchmark, we set $\alpha$ to be 1.5, as that parameter value gives the optimal performance. To consider integrability in this case, we take different values of $\alpha$ to determine how its integrability affects the performance. We will consider three different interaction ranges ($\ \alpha=0.0,1.5$, and $\infty$). The first $\alpha=0$ corresponds to all-to-all connectivity with equal strength, whereas $\alpha=\infty$ corresponds to nearest neighbour coupling only. Now $\alpha=1.5$ sits between these two extremes and corresponds to the long-range order parameter in recent ion-trap experiments [48, 8, 50, 51]. The training accuracy rates are summarized for these three parameter values in Fig. 5.
@@
Figure 5: Plot of the training accuracy rate of XX-Ising model for three different values of $\alpha=0.0$(blue), $1.5$(green), $\infty$(red). Here we have set $g/J_{0}=0.0$. The shadows displaced show the standard deviations obtained from 40 to 50 epochs. The inset table shows the unitary operators at the stroboscopic times, $J_{0}t=n\pi,(n+1/4)\pi,(n+1/2)\pi$ and $(n+1/3)\pi$ with $n\in\mathbb{N}_{\geq 0}$ and where $N$ is the number of sites and $\hat{P}_{X}$ the parity operator along the x-axis. The points and shadows are the mean value and the standard deviation taken from 40 to 50 epochs to reveal the stability of the optimization. (b) shows the $\alpha$-dependence with $g/J_{0}=0.5$. The blue and green dashed lines with dots plot the training and testing accuracy with $\alpha$ ranging from $\alpha=0.0$ to $10$ (the blue and green solid lines indicate the training and testing accuracy with $\alpha=\infty$).
@@
The best accuracy can be achieved for a middle range $\alpha$ value. Unlike the case of $\alpha=1.5$, the $\alpha=0$ and $\infty$ cases show periodical behavior. This reflects the periodic dynamics of the Hamiltonian with these particular values. It is useful to explore such mechanics more analytically. Our unitary operator can be decomposed as
@@
where the XX-interactions between each site commute with one another. In this case, the total unitary can be decomposed as a product of the simple oscillators with the angular frequencies $\omega_{lm}={J_{0}}/{|l-m|^{\alpha}}$. Thus, when $\alpha=1.5$, the oscillators are in different angular frequencies. This means the recurrence of dynamics is avoided, resulting in non-periodical dynamics.
Now, in the nearest-neighbour interaction case ($\alpha=\infty$), the oscillators may share the same angular frequency, introducing a symmetry. Our unitary operator may then be written as
@@
It is interesting also to explore the opposite limit of $\alpha=0$. Here, the Ising model also recovers the system symmetry, and the dynamics becomes solvable as
@@
with all oscillators having the same angular frequency $\omega_{lm}=J_{0}$. Thus, similar to $\alpha=\infty$, the dynamics is analytic at the times $J_{0}t=n\pi,(n+1/4)\pi,(n+1/2)\pi,(n+3/4)\pi$, where its performance reduces, as shown in Fig. 5. When the dynamics is periodical, unless the period is extremely long, it is likely that the dynamics can be fully supported by a small subspace of the Hilbert space. As shown in Fig. 5, the performance quickly recovers as $\alpha$ increases from zero, peaks around $\alpha=1$, and gradually decreases to the large $\alpha$ limit. This indicates that the non-integrability of the quantum system contributes to the performance of the quantum neural networks. Interestingly, the power to generate entanglement is not sufficient to estimate the QERC performance as we have seen in above with $\hat{U}^{(XX)}_{\alpha=\infty}$ at $J_{0}t=(n+1/4)\pi$ for instance.
This $\alpha$-dependence can be mitigated by applying a magnetic field in the $Z$-axis direction. In the ZZIsing model, the additional magnetic field in the $x$-axis direction breaks the U(1) symmetry. In contrast, in the XXIsing model, the additional field is necessary to resolve the integrability. Fig. 5 (b) shows the $\alpha$-dependency of the accuracy rate. The accuracy rate drops when $\alpha$ is large. The value of the transversal field strength is chosen to minimize the $\alpha$-dependency, which is $g/J_{0}=0.5$.
Integrability impacts the performance of the QERC to some degree. However, it does not give us clear criteria for the quantum reservoir. The ZZ-X model is an integrable system, though it does not hold U(1) symmetry due to its time evolution, which consists of two Hamiltonian components. This indicates that when the time-evolution has a “gate” structure the integrability may not work as an indicator. This is similar to the case of the Clifford gates $+$ T-gates. The Clifford gate circuit can be efficiently simulated, and does not give us high performance as a quantum reservoir. However, the one-depth of T-gates can be enough to achieve similar high performance to the Haar measure sampling. It may not be surprising that the integrability of the Hamiltonian system can significantly affect the performance of a quantum neural network. However, the integrability is much less important in the ZZ-X model, which is integrable with the analytical form for the output state. Integrability and periodicity are somewhat related in classical dynamics, and as we have seen above, the periodicity in the Hamiltonian dynamics prevents the states from evolving sufficiently for the one-layer classical neural network to distinguish their distributions. Non-integrability and non-periodicity of the system dynamics are not a necessary condition for a good quantum reservoir. However, they can still serve as an indication to provide good neural networks, as well as Haar random unitary does, which can be useful in designing quantum reservoirs.
@@ VI Information scrambling and sampling cost
In the previous section, we analyzed the properties of the quantum reservoirs with respect to their complex dynamics. Here, we focus on the other element of quantum computation that can be achieved by quantum reservoir, that is, the convergence to the solution through the quantum algorithm. So far, the quantum reservoirs we investigated require a linear classifier to complete classification tasks. If a quantum reservoir can converge the state well enough to make the linear classifier redundant, such a quantum reservoir would be ideal. The optimization of a quantum reservoir might be possible if we allow the quantum reservoir to be optimized. Our question here is whether such a computational mechanism can be achieved without optimization of the quantum reservoir.
To investigate this possibility, we look into the information scrambling and sampling cost for each quantum reservoir. The quantum dynamics provides a feature space for this computational model, and hence the dynamics complexity is important to deliver information scrambling necessary for the task.
A unitary matrix sampled on the Haar measure should give information scrambling in our benchmark. However, this process needs to be followed by the above computational mechanism, which is likely to be a clustering of states for classification problems. It is unclear if a quantum reservoir has any contribution to this process. If so, the quantum reservoir would help to work the linear classifier more efficiently. To investigate this, we need to estimate the sampling cost for the measurement.
@@
The amplitude distribution of the final state of the quantum reservoir was used as the output of the M-layer in our benchmark, however in the simulations, we used the amplitude distribution reconstructed from the data. Fig. 6 shows the behavior of the accuracy rates dependent on the sampling number. All the models in this comparison achieve similar accuracy rates with the theoretical amplitude distribution when the sampling number is large enough. It has often been claimed that the necessary sampling number grows exponentially with the number of qubits for quantum algorithms that require the probability amplitude of the output state. If this claim is correct, the sampling cost could be an obstacle for any experimental implementation. However, this conjecture does not apply to QERC, and we show the reason through the sampling analysis below.
To directly evaluate the reconstruction of a distribution, we can employ the Kullback-Leibler divergence (KLD) [64]. As we see in Fig. 11 (a) in Appendix B.2, the KLD scales polynomially for a fixed number of qubits for all four quantum reservoir models, and the KLD requires a larger sampling number to achieve the same KLD as the number of qubits increases. This might imply an exponential scaling cost for sampling as the number of qubits increases. However, the role of the M-layer is not to reconstruct the true amplitude distribution but to provide probability distributions for the linear classifier to classify them. Fig. 6 shows the convergence of the accuracy rate as the sampling number increases. In fact, it converges to the theoretical accuracy rate before the distributions are accurately reconstructed. Hence, in comparison to the cost of reconstructing the probability amplitude, the necessary sampling number is not high. The detailed analysis of the sampling cost for the linear classifier is in the Applendix B.2.
The remaining factor for the sampling cost is the scaling of the sampling number to the number of qubits in the quantum reservoir. For $N=12$, we obtain the same qualitative characteristics shown in Fig. 6 (see Fig. 11 (b-e) in the Appendix B.2 ). More importantly, the case of $N=12$ gives a higher accuracy rate for the same sampling number for all four models, and the accuracy rate converges to the theoretical value at the same sampling number for both numbers of qubits. These strongly indicate that the sampling cost in QERC is much lower than expected by the application of the standard sampling theory.
The analysis of the sampling cost also provides an insight into the differences among the quantum reservoirs. The ZZ-X model exhibits a different tendency in the small sampling number regime shown in Fig. 11, where it requires a much smaller sampling number than the other models require to achieve the same accuracy (see Appendix B.2). This can not only be a significant advantage in its implementation. But also be a hint to consider the trade-off between the information scrambling and the clustering of the output state to assist the classical linear classifier.
Information scrambling has to happen during the time evolution of the quantum reservoir. If this is only the role of the quantum reservoir, the unitary sampled on the Haar measure should give the best result. Our sampling cost analysis shows that the clustering of the state can happen with certain quantum reservoirs. In our case, the ZZ-X model achieves both information scrambling and partial clustering of the output states without any programming or optimization of the reservoir parameters. We investigate the participation ratio in the next section to examine this further.
@@ VII Participation Ratio (PR)
To evaluate the power of reservoir computation, the Lyapunov exponent and memory capacities such as $\tau$-capacity are widely used: the former being a measure of the chaotic dynamics and nonlinearlity in classical reservoirs [65], while the latter are used in classical and quantum reservoir computing [66, 30, 31, 67, 68, 69]. Memory capacity measures how much past information is stored in the reservoir, and unfortunately, such a measure is not appropriate for evaluating the quantum reservoirs used in QERC. Also, to see what is happening in the computational process through the QERC, we need to know what complexity is generated in the state as well as its degree. Hence, we employ the participation ratio (PR) used in condensed matter physics to measure the local and extended states [70], such as the Anderson localized and ergodic states, to evaluate the dynamics. For a quantum state $|\Psi\rangle=\sum_{l}c_{l}|l\rangle$, the PR is defined as,
@@
The PR of the output state can be highly dependent on the initial state, and so we considered the averaged PR (APR) over the output state set $\mathbb{D}$ defined by,
@@
This value shows how widely the state is spread over the computational basis on average, hence it tells us the representation capability that the quantum reservoir provides.
Now if the APR is high, an input state would evolve to an output state supported by many computational basis states. Hence, it is more likely that the dynamics yields a strong information scrambling. However, there is no guarantee that such a set of states is preferable to the classical linear classifier for classification. From the clustering point of view, all states for a certain class of images should have a distinct characteristic, and hence, APR cannot only describe the property of the quantum reservoir. The ideal distribution of the output states should have some clustering as well as exploiting the dimension of the Hilbert space. To see this, we introduce IAPR, which evaluates the overall usage of the dimensionality of the Hilbert space, as
${\mathrm{IAPR}}\left(\mathbb{D}\right)=\frac{{(\#\,\mathrm{states})}^{2}}{{\displaystyle\sum_{l}\left({\sum_{|\Psi\rangle\in\mathbb{D}}|\langle l|\Psi\rangle|^{2}}\right)^{2}}}.$ (11)
The IAPR is the APR for the distribution constructed from all the output states as
@@
If IAPR is large, it indicates that the dimensionality of the Hilbert space is exploited for the set of input states. Now, we consider the difference, $\Delta\mathrm{PR}=\mathrm{IAPR}-\mathrm{APR}$. If this difference is large, each state is supported by a relatively small number of basis states, whereas the set of states requires the entire dimension of the Hilbert space. Hence, this difference can be considered an indicator for clustering the output states.
@@
Before analyzing the PRs, we need to consider the effect of the bias within the set of input states. These PR values depend on the trend and bias within the set of input states. To evaluate the quantum reservoirs independently of these trends, we introduce two random input sets. First, we consider uniformly distributed points in the $2^{N}$-dimensional space. We generate $10,000$ points randomly and use the same encoder as we did in the benchmark (described in Ref. [21] and the Appendix A.1). However, the completely random input set is not so realistic, as the input data set generally has some bias. Hence, we also consider a random data set generated by isotropic Gaussian blobs in the high-dimensional square with length $-10$ to $10$ (see the sketch in Fig. 7 (b-1)). Here we randomly generate $10,000$ points in the high-dimensional space. These points form $10$ isotropic Gaussian blobs (clusters) where the standard deviation of the clusters is $1.2$. These points are mapped to the angles following the encoding process mentioned in Ref. [21] (and Appendix A.1). We also considered the MNIST data set to evaluate the actual practical case (see (c-1) in Fig. 7).
@@
Figure 7: Plot of the APR vs. $\Delta\mathrm{PR}$ for ZZX-Ising model using ten qubits. The red and green gradation points show the result with different transverse magnetic strengths $0\leq g/J_{0}\leq 5$, where we used normal and randomized encoders in red and green, respectively. Here, we set $J_{0}t=3.5$ with $\alpha=1.5$. The blue dots show the changes in APR and IAPR over time. The red and blue triangle markers show the results of the Haar and SRC reservoirs. Now (a) and (b) show the results of sets where $10,000$ initial states were randomly generated. For (a), all states were created by angles $\theta_{l}$ and $\phi_{l}$ that are uniformly randomly generated. Next, (b) illustrates the result for the initial state set generated by isotopic Gaussian blobs, while (c) depicts the results for the MNIST dataset using 60,000 images.
@@
We show the APR and $\Delta\mathrm{PR}$ values for the ZZ-Ising model with the benchmark against two random reservoirs: the Haar measure sampling and the single rotation random circuit (SRC). These are given by the red and blue triangles in Fig. 7, respectively. For the ZZ-Ising situation, the difference between the randomization of the input state appears in the difference between the red (no randomization) and green (randomization) dots. As the magnetic field strength $g$ increases, both the APR and IAPR decrease, as shown with the red dots. This agrees with the lower accuracy without randomization for the high $g$ regime. For all three data sets, the ZZ-Ising model gets close to the value given by the Haar sampling, and hence by taking optimal parameter values, the behaviour of this reservoir is expected to be similar to that of Haar sampling. However, the optimal value for $g$ is not where the APR is the largest. To see this more clearly, we need to investigate the PR values for the ZZ-X model.
Figure 8: Normalized kernel functions for QORC and RFF: We randomly selected 512 images ($X_{i=1,2,\ldots,512}$) from the MNIST data set (training set) and computed normalized kernel functions for QORC and RFF. The upper figures are kernel functions for QORC, and the lower figures are kernel functions for RFF. The two axes correspond to the image indices, and the color represents the kernel value. The number of features for both models is $2^{N}$, where $N$ is from 8 to 12.
The APR results show that the states are highly spread out in Haar, utilizing a vast number of computational bases. This suggests that distinguishing between the states of different images is challenging. However, considering $\Delta$PR, the PR per image is never the entire space, i.e., $2^{N}$, approximately half of it. When considering all images collectively, the total space is covered. This implies that images within the same class significantly overlap the computational bases used, whereas images from different classes have fewer common computational bases. To further clarify this point, we examine the kernel matrix, which is widely used in machine learning analysis of feature mappings [72]. This kernel matrix is the inner product of feature mapping vectors for two inputs, indicating the similarity between data points in feature space. Intuitively, the kernel values are high for images within the same class and lower for those in different classes.
Specifically, we randomly selected 512 images from the training dataset, with each class having approximately 50 images on average. The probability distributions for these images are obtained using Haar-random unitary matrices as a reservoir. Considering our QERC post-processing, we compute the normalized kernel matrix for the outputs processed with Eq. 15 in the Appendix A and compare it with the classical random feature model, Random Fourier Features (RFF) [73, 74]. Now, the normalized kernel matrix is given by,
$K_{\left(i,j\right)}(x_{i},x_{j})=\frac{q\left(x_{i}\right)^{T}\cdot q\left(x_{j}\right)}{\sqrt{q\left(x_{i}\right)^{T}\cdot q\left(x_{i}\right)}\sqrt{q\left(x_{j}\right)^{T}\cdot q\left(x_{j}\right)}},$ (13)
where $x_{i,j}$ are images and $q_{i,k}$ are corresponding feature map vectors. Note that, in the RFF calculation, we prepared $2^{N}$ features, where $N$ is the number of qubits we used.
The results are shown in Fig. 8. High similarity is observed between data points within the same class, whereas similarity is significantly lower between different classes. Furthermore, as the number of qubits increases, the similarity between different classes decreases, leading to better class separation. This suggests that our model possesses strong discriminative power in high-dimensional spaces. Additionally, compared with RFF, our model exhibits lower similarity between different classes, enabling more effective discrimination of distinct distributions. At the same time, high similarity is maintained within the same class, ensuring appropriate classification while mitigating overfitting risks.
@@
The PR values for the ZZ-X model are shown in Fig. 9, with an interesting characteristic being that there are points above the Haar sampling in terms of the APR in the case of the ZZ-X model. Even more interestingly, the optimal point for the QERC performance does not appear near the Haar sampling. The blue dots in these figures show the dependency of the PRs on the single-rotation angle. The optimal point for $\theta_{x}$ in Fig. 4 (c) is near $\pi/8$, meaning the optimal point for this model clearly differs from the point given by Haar sampling. Further the ZZ-X model optimal point is characterized by a lower APR and higher IAPR. For the cases with a lower APR and a higher IAPR, each output distribution function does not broadly spread on the computational basis. This supports the idea that the ZZ-X model utilizes the Hilbert space as much as other models to support the set of output states, although the distribution of each output state is not so flat. This clustering effect assists the classical linear classifier distinguish them, reducing the sampling cost. This PR ratio and the high performance of the ZZ-X model provide evidence that the ZZ-X model realizes the clustering of the output states, meaning that the process to converge to the solution of the task has already started in the time evolution of the quantum reservoir.